# NSP: A Neuro-Symbolic Natural Language Navigational Planner

### v1

@@ Abstract
Path planners that can interpret free-form natural language instructions hold promise to automate a wide range of robotics applications. These planners simplify user interactions and enable intuitive control over complex semi-autonomous systems. While existing symbolic approaches offer guarantees on the correctness and efficiency, they struggle to parse free-form natural language inputs. Conversely, neural approaches based on pre-trained Large Language Models (LLMs) can manage natural language inputs but lack performance guarantees. In this paper, we propose a neuro-symbolic framework for path planning from natural language inputs called NSP. The framework leverages the neural reasoning abilities of LLMs to i) craft symbolic representations of the environment and ii) a symbolic path planning algorithm. Next, a solution to the path planning problem is obtained by executing the algorithm on the environment representation. The framework uses a feedback loop from the symbolic execution environment to the neural generation process to self-correct syntax errors and satisfy execution time constraints. We evaluate our neuro-symbolic approach using a benchmark suite with 1500 path-planning problems. The experimental evaluation shows that our neuro-symbolic approach produces 90.1% valid paths that are on average 19-77% shorter than state-of-the-art neural approaches.
@@ III Problem Formulation
There is a house with Room1, Room2, Room3, and Room4. Room1 is connected to Room2 and Room3, Room2 is connected to Room3 and Room4, and Room3 is connected to Room4. Start in Room1 and move to Room2. Do not pass through Room4.” The expected output from this problem instance would be “[Room1, Room3, Room2]".
There is a house with Room1, Room2, Room3, and Room4. Room1 is connected to Room2 with a weight of $3$ and Room4 with a weight of $5$, Room2 is connected to Room4 with a weight of $3$ and Room3 with a weight of $2$, and Room3 is connected to Room4 with a weight of $1$. Find a path that begins and ends in Room1 that passes through each room at least once.” The expected output from this problem instance would be “[Room1, Room4, Room3, Room2, Room1]".
@@ IV-B Neuro-Symbolic Planning & Feedback
if error_message: {prompt} += “An error occurred with the previous response: {function_code} The error message was: {error_message} Please correct the response." *if code timed out, error_message = "The code took too long to execute. Increase the time efficiency or approximate the solution."
@@ IV-C Case Study on a Small Example
The case study considers the following problem description $\mathcal{D}$ “I have a house with the following rooms: Room1, Room2, Room3, Room4, Room5, Room6, Room7, Room8, Room9, and Room10. Room1 is connected to Room2 with a distance of 8, and Room10 with a distance of 8. Room2 is connected to Room6 with a distance of 7, Room10 with a distance of 5, and Room8 with a distance of 4. Room3 is connected to Room10 with a distance of 3, and Room7 with a distance of 6. Room6 is connected to Room8 with a distance of 6. Room7 is connected to Room9 with a distance of 7. Room8 is connected to Room10 with a distance of 2. Start in Room6 and go to Room9 without passing through Room5, and Room4. Avoid moving from Room5 directly into Room4, Room5 directly into Room8, Room8 directly into Room1, and Room1 directly into Room4."

### v2

@@ Abstract
Abstract Path planners that can interpret free-form natural language instructions hold promise to automate a wide range of robotics applications. These planners simplify user interactions and enable intuitive control over complex semi-autonomous systems. While existing symbolic approaches offer guarantees on the correctness and efficiency, they struggle to parse free-form natural language inputs. Conversely, neural approaches based on pre-trained Large Language Models (LLMs) can manage natural language inputs but lack performance guarantees. In this paper, we propose a neuro-symbolic framework for path planning from natural language inputs called NSP. The framework leverages the neural reasoning abilities of LLMs to i) craft symbolic representations of the environment and ii) a symbolic path planning algorithm. Next, a solution to the path planning problem is obtained by executing the algorithm on the environment representation. The framework uses a feedback loop from the symbolic execution environment to the neural generation process to self-correct syntax errors and satisfy execution time constraints. We evaluate our neuro-symbolic approach using a benchmark suite with 1500 path-planning problems. The experimental evaluation shows that our neuro-symbolic approach produces 90.1% valid paths that are on average 19-77% shorter than state-of-the-art neural approaches.
@@ II   Related Work
Large language models (LLMs), such as those from Anthropic and OpenAI, have demonstrated remarkable capabilities in generating human-like text and solving complex tasks, including natural language processing, code generation, and planning. However, these models are known to produce hallucinations, i.e., factually incorrect or contextually inappropriate outputs, which poses a significant risk particularly in safety-critical applications like autonomous systems and planning tasks. Recent work [27, 28] has addressed this issue by combining inductive reasoning from LLMs with deductive verification methods such as satisfiability solving. This approach uses formal methods to detect and correct erroneous outputs from LLMs, iteratively guiding them towards producing accurate and verified solutions. By integrating counterexample-guided inductive synthesis, the model is refined through feedback from deductive solvers until a valid solution is obtained. This methodology significantly enhances the trustworthiness and reliability of LLMs in applications where correctness is crucial, such as planning in autonomous systems.
@@ III   Problem Formulation
There is a house with Room1, Room2, Room3, and Room4. Room1 is connected to Room2 and Room3, Room2 is connected to Room3 and Room4, and Room3 is connected to Room4. Start in Room1 and move to Room2. Do not pass through Room4.” The expected output from this problem instance would be “[Room1, Room3, Room2]”.
There is a house with Room1, Room2, Room3, and Room4. Room1 is connected to Room2 with a weight of $3$ and Room4 with a weight of $5$, Room2 is connected to Room4 with a weight of $3$ and Room3 with a weight of $2$, and Room3 is connected to Room4 with a weight of $1$. Find a path that begins and ends in Room1 that passes through each room at least once.” The expected output from this problem instance would be “[Room1, Room4, Room3, Room2, Room1]”.
@@ IV-B   Neuro-Symbolic Planning & Feedback
if error_message: {prompt} += “An error occurred with the previous response: {function_code} The error message was: {error_message} Please correct the response.” *if code timed out, error_message = ”The code took too long to execute. Increase the time efficiency or approximate the solution.”
@@ IV-C   Case Study on a Small Example
The case study considers the following problem description $\mathcal{D}$ “I have a house with the following rooms: Room1, Room2, Room3, Room4, Room5, Room6, Room7, Room8, Room9, and Room10. Room1 is connected to Room2 with a distance of 8, and Room10 with a distance of 8. Room2 is connected to Room6 with a distance of 7, Room10 with a distance of 5, and Room8 with a distance of 4. Room3 is connected to Room10 with a distance of 3, and Room7 with a distance of 6. Room6 is connected to Room8 with a distance of 6. Room7 is connected to Room9 with a distance of 7. Room8 is connected to Room10 with a distance of 2. Start in Room6 and go to Room9 without passing through Room5, and Room4. Avoid moving from Room5 directly into Room4, Room5 directly into Room8, Room8 directly into Room1, and Room1 directly into Room4.”